OpenStudy (anonymous):

evaluate the following apply limit lim theta->0 sin(under root 2theta)/(underroot2theta)

OpenStudy (anonymous):

hope u can undrstand the question

OpenStudy (anonymous):

1

OpenStudy (anonymous):

i need the solution

OpenStudy (anonymous):

:(

OpenStudy (angela210793):

\[\lim_{\theta \rightarrow 0} \sin \sqrt{\theta}\div \sqrt{\theta}=1\]

OpenStudy (anonymous):

let sqrt(2theta)=t then u get sint / t t ->0

OpenStudy (angela210793):

lim X-->0 of sinx/x=1 tht's smth tht is given and u have to remember ...

OpenStudy (anonymous):

1 is the limit

OpenStudy (anonymous):

what you need to know to do this is that \[lim_{x->0}\frac{sin(x)}{x}=1\]

OpenStudy (anonymous):

of course that limit is no different than \[lim_{x->0}\frac{sin(2x)}{2x}=1\]

OpenStudy (anonymous):

and the square root of one is one.

OpenStudy (anonymous):

then again maybe it was \[lim_{x->0} \frac{sin(\sqrt{x})}{\sqrt{x}}\] which is of course the same

OpenStudy (anonymous):

or even \[lim_{x->0}\frac{sin(2\sqrt{x})}{2\sqrt{x}}\]

OpenStudy (anonymous):

are you convinced of this or no?

OpenStudy (angela210793):

No1 would have explained it better :):) G j Satellite :):)

OpenStudy (anonymous):

you know now that i think about it maybe this is a trick

OpenStudy (anonymous):

and you are supposed to say that the limit does not exist

OpenStudy (anonymous):

because the domain of thing beast is \[x\geq 0\] and therefore you cannot approach 0 from the left

OpenStudy (anonymous):

now that i think of it...i think satellites right

OpenStudy (anonymous):

so perhaps that was the point of this question

OpenStudy (angela210793):

W8.....explain me again plsss

OpenStudy (anonymous):

the lhl wouldnt exist

OpenStudy (anonymous):

see the right hand limit is 1, and that part is fine

OpenStudy (anonymous):

unless it says\[lim_{x->0^+}\frac{sin(\sqrt{x})}{\sqrt{x}}\]

OpenStudy (angela210793):

oh..i got it!!

OpenStudy (anonymous):

but from the left side, if u put 0-h for theta, ud get the root of a negative number, which doesnt exist, for real

OpenStudy (anonymous):

cannot take a limit because the function does not exist to the left. on account of the square root.

OpenStudy (anonymous):

what him1618 said

OpenStudy (anonymous):

satellite keeps on proving how amazing he is

OpenStudy (anonymous):

yeah right

OpenStudy (anonymous):

but i will thank you for the compliment

OpenStudy (angela210793):

Yea.yea.....u 2 r right....Thnx :) (Satellite is a he?) oO

OpenStudy (anonymous):

does it say limit from the left or just limit?

OpenStudy (anonymous):

r u a he satellite?

OpenStudy (anonymous):

i'll never tell

OpenStudy (angela210793):

I've always thought he/she is a she...

OpenStudy (anonymous):

beauty of the world wide interweb is you can be anyone. a dog even

OpenStudy (anonymous):

im sure its a she....

OpenStudy (angela210793):

I'm pretty sure is a she too :P so whtever ur....for us ur a she :P

OpenStudy (anonymous):

amazing then..some gorgeous lady supporting a 180 iq

OpenStudy (angela210793):

whose IQ is 180? urs?

OpenStudy (anonymous):

hehe thanx every 1 satellite,him1618 and angela:) thanx a lot

OpenStudy (anonymous):

satellites i think..lol..m talkin abt her

OpenStudy (anonymous):

welcome mimi

OpenStudy (angela210793):

Haha lol :P:P :) uw :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

btw angela i first fell in love on 21st july...:)

OpenStudy (anonymous):

:)

OpenStudy (angela210793):

wht year?

OpenStudy (anonymous):

2007

OpenStudy (anonymous):

we were still 14 then...

OpenStudy (angela210793):

u remember the date? Interesting...I probably wouldn't :P:P

OpenStudy (anonymous):

i remember a lot...lol...no use though..no response....

OpenStudy (angela210793):

I remember only dates tht have to do with history lessons :P:P And I'm sorry :S

OpenStudy (anonymous):

sweet of u...i never remembered history..except perhaps the knights templar and stuff...

OpenStudy (angela210793):

I remember dates such as when was Magna Carta was signed,or when did the Roman Empire split in 2...these things :P..nothing abt albanian history though,...only independence day :P:P

OpenStudy (anonymous):

hahaha....

OpenStudy (angela210793):

:D:D ^_^

OpenStudy (mathteacher1729):

Here is a great little visualization and explanation of the basic limit \(\lim_{x\to 0}\frac{\sin(x)}{x}\). http://www.ies.co.jp/math/java/calc/LimSinX/LimSinX.html Hope this helps. :)